A ball is launched from the lab table, with an initial velocity of 2.5 m/s. If the table is 0.7 m high, how far from the table does it land? How long does it take to land?
1
Expert's answer
2020-12-17T09:01:17-0500

How long it will take to land:


"t=\sqrt{\frac{2h}{g}}=\sqrt{\frac{2\cdot0.7}{9.8}}=0.378\text{ s}."

How far from the table it will land:


"R=vt=2.5\cdot0.378=0.945\text{ m}."